# Periodic Korteweg-de Vries soliton potentials generate magnetic field strength with excellent quasisymmetry

### v2

@@ Acknowledgments
We conclude by thanking P. Helander, B. Khesin, J.W. Burby, M. Wheeler, J. Meiss, N. Duignan, V. Duarte, H. Weitzner, E. Rodriguez, F.P. Diaz and M. Landreman for helpful discussions and suggestions. This research was supported by a grant from the Simons Foundation/SFARI (560651, AB) and DoE Grant No. DE-AC02-09CH11466.
@@ Appendix A The QS condition $\bm{u}\cdot\bm{\nabla}B=0$ in Clebsch variables
where $\iota,\psi$ denotes the rotational transform and the poloidal flux. It is convenient to define a flux coordinate [48] $(\psi,\theta,\phi)$ such that the Jacobian $1/\sqrt{g}=\bm{\nabla}\psi\times\bm{\nabla}\theta\cdot\bm{\nabla}\phi=F/R^{2}$. In the flux coordinates, we have
@@
A similar analysis can be carried out in standard Boozer coordinates $(\psi,\vartheta_{B},\phi_{B})$ [27] where
@@ Appendix B Derivation of the KdV equation for vacuum quasisymmetric B in slab geometry
We shall now present a detailed derivation of the KdV equation in slab geometry. There are several steps in the calculation and we shall present them in the following subsections. First, we show the consistency of (24c) with the two-term form of QS. Second, we derive the relation between $(\psi,\alpha,\ell)$ and the Cartesian coordinates. Third, we derive the expressions for $H$. Finally, we expand the GB and the KBC equations order by order.
@@ B.1 Equivalence of (24c) and the two-term form of QS
Thus, for vacuum fields $B$ satisfies a travelling-wave-like solution [47] in $(\Phi,\alpha)$ of the form [47]
@@ C.2 Non-orthogonal Boozer Coordinates
The Jacobian of the $(\psi_{s},\alpha_{s},\Phi_{s})$ system is $J^{-1}=B^{2}_{s}$ The contravariant vectors [48] are easily seen to be
@@ Appendix D Derivation of $\mathcal{D}(\psi)$ in equation (12)
in terms of the roots $B_{M}(\psi),B_{m}(\psi),B_{X}(\psi)$, $\iota(\psi)$, the rotational transform, and the quasihelicity $N$ [11, 27].
@@
Note that $G(\psi),I(\psi)$ are flux functions related to the poloidal and toroidal currents. [11] Since $\bm{B}\cdot\bm{\nabla}=B\partial_{\ell}$ in $(\psi,\alpha,\ell)$, (122) takes the form
@@
Finally, we shall need some basic identities satisfied by elliptic functions (see Appendix A of reference [28]). In particular, we shall use
@@
where dn$(z,m)$, cn$(z,m)$ are standard Jacobi elliptic functions, and K$(m)$, E$(m)$ are the complete elliptic integrals of the first and second kind.[30]
@@ Appendix E Additional configurations that satisfy the integrated KdV equation (12)
The precise QA configuration with larger negative shear in Fig. 21 was obtained by a series of optimizations starting from the precise QA.[4] Specifically, we start from the same objective function used to optimize the precise QA, and the precise QA configuration itself. The mean shear $\bar{\hat{s}}$ was extracted from the precise QA configuration, and a new least-squares term $(\bar{\hat{s}}-\bar{\hat{s}}_{*})^{2}$ was added to the objective function, where the target value for mean shear $\bar{\hat{s}}_{*}$ is taken to be slightly different from the extracted value. This new objective is then optimized locally, using the precise QA as the initial state. This ensures that the additional least-squares term is almost zero, so the precise QA is approximately an optimum to the modified objective function. Hence, the new local optimum can be expected to be close to the initial configuration. This process is then repeated starting from the slightly modified optimum, each step generating a similar equilibrium with a slightly different mean shear. Mean shear is here defined as $b/\bar{\iota}$, where $b$ is the result of fitting a first-order polynomial $a+bs$ to the iota profile, where $s$ is the normalized toroidal flux; $\bar{\iota}$ is the rotational transform averaged over all radii. The mean shear in the original precise QA is about $-0.018$, and we incremented the target negative shear in steps of $0.005$.
Figure 20: The value of $(dB/d\ell)^{2}$ on the outermost flux surface in the precise QH [4] as a function of $B$, alongside a cubic polynomial fit (a), and the errors of quadratic, cubic, and quartic fits on each flux surface (b). Panel (c) shows the behavior of the roots in the precise QH.
@@
Figs. 23 and 24 show equilibria that were not derived from the precise QA/QH. Instead, they were only optimized for quasisymmetry on the outermost flux surface, with the quasisymmetry degrading as the axis is approached. For both cases, the ratio of symmetry-breaking modes to quasisymmetric modes near the axis is $\sim 10^{-3}$. While this may seem a small value, the resulting ripple creates enough noise in the $(dB/d\ell)^{2}$ vs. $B$ plot to render it meaningless, and the error for all fits is $\sim 1$ in that region. The symmetry-breaking modes were filtered out when calculating the roots shown in panels (c). Note that the equilibrium shown in Fig. 23 is the same one used in Fig. 2 in the main text.

### v3

@@ Acknowledgements.
We thank P. Helander, B. Khesin, J.W. Burby, M. Wheeler, J. Meiss, N. Duignan, V. Duarte, H. Weitzner, E. Rodriguez, F.P. Diaz and M. Landreman for helpful discussions and suggestions. This research was supported by a grant from the Simons Foundation/SFARI (560651, AB) and DoE Grant No. DE-AC02-09CH11466.
@@ Appendix A The QS condition $\bm{u}\cdot\bm{\nabla}B=0$ in Clebsch variables
where $\iota,\psi$ denotes the rotational transform and the poloidal flux. It is convenient to define a flux coordinate (D’haeseleer et al., 2012) $(\psi,\theta,\phi)$ such that the Jacobian $1/\sqrt{g}=\bm{\nabla}\psi\times\bm{\nabla}\theta\cdot\bm{\nabla}\phi=F/R^{2}$. In the flux coordinates, we have
@@
A similar analysis can be carried out in standard Boozer coordinates $(\psi,\vartheta_{B},\phi_{B})$ Landreman and Sengupta (2018) where
@@ Appendix B Derivation of the KdV equation for vacuum quasisymmetric B in slab geometry
We now present a detailed derivation of the KdV equation in slab geometry. There are several steps in the calculation and we shall present them in the following subsections. First, we show the consistency of (24c) with the two-term form of QS. Second, we derive the relation between $(\psi,\alpha,\ell)$ and the Cartesian coordinates. Third, we derive the expressions for $H$. Finally, we expand the GB and the KBC equations order by order.
@@ B.1 Equivalence of (24c) and the two-term form of QS
Thus, for vacuum fields $B$ satisfies a travelling-wave-like solution (Sengupta et al., 2021) in $(\Phi,\alpha)$ of the form Sengupta et al. (2021)
@@ C.2 Non-orthogonal Boozer Coordinates
The Jacobian of the $(\psi_{s},\alpha_{s},\Phi_{s})$ system is $J^{-1}=B^{2}_{s}$ The contravariant vectors (D’haeseleer et al., 2012) are easily seen to be
@@ Appendix D Derivation of $\mathcal{D}(\psi)$ in equation (12)
in terms of the roots $B_{M}(\psi),B_{m}(\psi),B_{X}(\psi)$, $\iota(\psi)$, the rotational transform, and the quasihelicity $N$ Helander (2014); Landreman and Sengupta (2018).
@@
Note that $G(\psi),I(\psi)$ are flux functions related to the poloidal and toroidal currents. Helander (2014) Since $\bm{B}\cdot\bm{\nabla}=B\partial_{\ell}$ in $(\psi,\alpha,\ell)$, (122) takes the form
@@
Finally, we shall need some basic identities satisfied by elliptic functions (see Appendix A of (Kamchatnov, 2000)). In particular, we shall use
@@
where dn$(z,m)$, cn$(z,m)$ are standard Jacobi elliptic functions, and K$(m)$, E$(m)$ are the complete elliptic integrals of the first and second kind.Kamchatnov (2000)
@@ Appendix E Additional configurations that satisfy the integrated KdV equation (12)
The precise QA configuration with larger negative shear in Fig. 21 was obtained by a series of optimizations starting from the precise QA.Landreman and Paul (2022) Specifically, we start from the same objective function used to optimize the precise QA, and the precise QA configuration itself. The mean shear $\bar{\hat{s}}$ was extracted from the precise QA configuration, and a new least-squares term $(\bar{\hat{s}}-\bar{\hat{s}}_{*})^{2}$ was added to the objective function, where the target value for mean shear $\bar{\hat{s}}_{*}$ is taken to be slightly different from the extracted value. This new objective is then optimized locally, using the precise QA as the initial state. This ensures that the additional least-squares term is almost zero, so the precise QA is approximately an optimum to the modified objective function. Hence, the new local optimum can be expected to be close to the initial configuration. This process is then repeated starting from the slightly modified optimum, each step generating a similar equilibrium with a slightly different mean shear. Mean shear is here defined as $b/\bar{\iota}$, where $b$ is the result of fitting a first-order polynomial $a+bs$ to the iota profile, where $s$ is the normalized toroidal flux; $\bar{\iota}$ is the rotational transform averaged over all radii. The mean shear in the original precise QA is about $-0.018$, and we incremented the target negative shear in steps of $0.005$.
Figure 20: The value of $(dB/d\ell)^{2}$ on the outermost flux surface in the precise QH Landreman and Paul (2022) as a function of $B$, alongside a cubic polynomial fit (a), and the errors of quadratic, cubic, and quartic fits on each flux surface (b). Panel (c) shows the behavior of the roots in the precise QH.
@@
We show equilibria that were not derived from the precise QA/QH in Figures 23 and 24. Instead, they were only optimized for quasisymmetry on the outermost flux surface, with the quasisymmetry degrading as the axis is approached. For both cases, the ratio of symmetry-breaking modes to quasisymmetric modes near the axis is $\sim 10^{-3}$. While this may seem a small value, the resulting ripple creates enough noise in the $(dB/d\ell)^{2}$ vs. $B$ plot to render it meaningless, and the error for all fits is $\sim 1$ in that region. The symmetry-breaking modes were filtered out when calculating the roots shown in panels (c). Note that the equilibrium shown in Fig. 23 is the same one used in Fig. 2 in the main text.